# Simulating chiral spin liquids with fermionic Projected Entangled Paired States

### v1

@@ I Introduction
Chiral spin liquids are exotic states of matter characterized by the absence of magnetic ordering while breaking time-reversal ($T$) and parity ($P$) symmetries [1]. They also exhibit long-range topological order [2]. They have been encountered in several quantum spin models with $\rm SU(2)$ [3] or higher $\rm SU(N)$ [4] symmetry in the presence of a chiral term breaking explicitly $T$ and $P$. In some cases, $P$ and $T$ can be broken spontaneously [5, 6].
Tensor networks like Projected Entangled Pair States (PEPS) [7] are well suited to the investigation of spin liquids. Topological orders can be encoded naturally by imposing virtual gauge symmetries [8, 9]. In addition, chiral forms of PEPS can describe CSL [10, 11] and be used as an efficient variational scheme to attack frustrated quantum spin models hosting CSL phases [12, 13]. On the other hand, infinite PEPS (iPEPS) is an ideal tool as it defines states in the thermodynamic limit directly, avoiding finite size extrapolations.
@@ II.1 Parton ansatze
where $\langle ij\rangle$ ($\langle\langle ik\rangle\rangle$) denotes nearest (next-nearest) neighbor bonds. We fix $t_{2}=0.5t_{1}$, $\chi_{ij}=\pm 1$ to ensure a $\pi$ flux through every square plaquette and choose the complex phases $\theta_{ik}$ to obtain a $\pi/2$ flux in all triangles. In this paper we choose the gauge used in Ref. [24] such that the unit-cell can be chosen as two nearest neighbour sites along the $x$ direction.
The exact many-body parton wavefunction $|\varphi\rangle$ is a Gutzwiller projected Slater-determinant
@@
It is known that the resultant exact parton state is a $\rm SU(2)_{1}$ CSL which is equivalent to the $\nu=1/2$ bosonic Laughlin state. On a torus, such CSL has two-fold topological degeneracy where the degenerate states can be constructed by imposing different boundary conditions on the parton wavefunctions [25]. On a cylinder, the CSL hosts chiral gapless edge states predicted by $\rm SU(2)_{1}$ Wess-Zumino-Witten (WZW) CFT.
@@ III.2 Spin-spin correlations
Figure 2: Logarithmic plot of the magnitude of spin-spin correlations as a function of distance for the exact and projected fPEPS, for the $L=18$ cluster with periodic boundary conditions. Symbols are ommited if the obtained value is lesser than one standard deviation of error. The expected correlations of the exact CSL at $r>5$ (in the thermodynamic limit) are shown as a dotted line, extending the exponential behavior.
@@ III.3 Topological properties: ground state degeneracy
$|\Psi^{\text{PBC-PBC}}\rangle&=\frac{|\Psi^{\text{PBC-APBC}}\rangle+|\Psi^{\text{APBC-PBC}}\rangle}{\sqrt{2}},\\ |\Psi^{\text{APBC-APBC}}\rangle&=\frac{|\Psi^{\text{PBC-APBC}}\rangle-|\Psi^{\text{APBC-PBC}}\rangle}{\sqrt{2}},\end{aligned}$ (8)
with a very good accuracy whenever $L\geq 18$. These results substantiate that fPEPS, even at finite bond dimensions, can accurately capture the correct topological degeneracy of CSL states.
@@ IV Simulation of the chiral $J_{1}-J_{2}-J_{\chi}$ model
Note that this model has been investigated in Ref. [12], where the sum of four triangular $J_{\chi}$ terms inside a plaquette has been equivalently written as a spin cyclic permutation $i(P_{ijkl}-P_{ijkl}^{-1})$ term. Ref. [12] performed a variational study of this model with bosonic iPEPS, and found a regime of $\rm SU(2)_{1}$ CSL in the phase diagram. The optimized bosonic iPEPS provides good variational energy and correct level counting in the ES predicted by $\rm SU(2)_{1}$ WZW CFT. However, there exists a redundant chiral branch in the odd (semion) sector [12, 36] which contradicts both the theoretical prediction and recent numerical results obtained from DMRG on finite cylinders [22, 23]. We emphasize that such artificial replication of chiral branches in bosonic iPEPS is quite general and is also found in the cases of $\rm SU(N)$ and non-Abelian CSLs [19, 20, 4, 13].
@@ IV.1 Variational energy
We choose the parameters of the spin model as $J_{1}=2\cos(0.06\pi)\cos(0.14\pi)$, $J_{2}=2\cos(0.06\pi)\sin(0.14\pi)$, $J_{\chi}=4\sin(0.06\pi)$ which has been considered in Refs. [12, 34] and is known to be deep inside the CSL phase with the ground state energy being $E\approx-1$. We import the $\rm SU(2)$ bosonic iPEPS method there and compute variational energies as a reference. For the bosonic iPEPS, the tensor has $\rm SU(2)$ symmetry and the unit-cell size is chosen to be one such that each tensor only has one physical leg. The results for virtual spaces $V=0\oplus 1/2$ ($D=3$) and $V=0\oplus 0\oplus 1/2$ ($D=4$) are given in Fig. 4. As we extrapolate to the infinite $\chi$ limit the energies are around $-0.99$.
@@ V Conclusion and outlook
In this work, we investigated topological properties of the projected fPEPS ansatz for CSL state. By performing a VMC analysis on the initial parton state, we demonstrated that fPEPS at finite bond dimension can accurately capture the topological GS degeneracy and spin-spin correlations of CSLs. We expect similar behaviours for the modular S and T matrices as indicated by the high quality of overlap fidelity.
@@
Our fPEPS ansatz can further be used to study fermionic Hofstadter-Hubbard model [38] where both Chern insulating phase and CSLs can be tuned by strength of Hubbard interaction (e.g. Ref. [39]). In that case, our Gaussian fPEPS ansatze with partially projected doublons are expected to be good initial variational ansatze at finite Hubbard interaction.

### v2

@@ I Introduction
Commonly, a quantum spin liquid refers to a quantum state that does not break spontaneously the system symmetries, neither the space group symmetry nor the global continuous (e.g. SU(2)) symmetry, if any of those is present in the Hamiltonian. In particular, spin liquids show absence of magnetic ordering down to zero temperature. Chiral spin liquids (CSLs) are exotic states of matter characterized, in addition to their spin liquid character, by the breaking of time-reversal ($T$) and parity ($P$) symmetries [1]. They also exhibit long-range topological order [2]. They have been encountered in several quantum spin models with $\rm SU(2)$ [3, 4, 5, 6, 7, 8, 9, 10] or higher $\rm SU(N)$ [11, 12, 13, 14, 15, 16, 17, 18] symmetry in the presence of a chiral term breaking explicitly $T$ and $P$. In some cases, $P$ and $T$ can be broken spontaneously, as demonstrated in both Hubbard/Heisenberg like models [19, 20, 21, 22, 23, 24, 25, 26] and Kitaev models [27] on lattices with odd plaquettes [28, 29, 30, 31, 32, 33, 34, 35]. Parent hamiltonians have also been devised for CSLs [36, 37, 38].
Experimental efforts to identify compounds with CSL ground states are ongoing, with prominent attention devoted to spin-orbit coupled materials with Kitaev-like interaction under external magnetic fields [39, 40, 41]. Additionally, in Moiré heterostructures, the layer degree of freedom regarded as a pseudo-spin facilitates realization of pseudo-chiral spin liquids [42]. Chiral spin liquids have also been proposed in cold atom platforms using Floquet driving [43] or Rydberg arrays [44, 45, 46, 47] where the chirality of edge modes can be detected experimentally. CSLs relevant for Rydberg atom models on the honeycomb lattice have recently been classified via a projective symmetry group approach [48].
As CSLs typically emerge in strongly correlated systems, numerical tools for determining phase diagrams of generic Hamiltonians and characterizing CSLs are desired. Tensor networks like Projected Entangled Pair States (PEPS) [49] are well suited to the investigation of spin liquids. Topological orders can be encoded naturally by imposing virtual gauge symmetries [50, 51]. In addition, chiral forms of PEPS can describe CSL [52, 53] and be used as an efficient variational scheme to attack frustrated quantum spin models hosting CSL phases [54, 55]. On the other hand, infinite PEPS (iPEPS) is an ideal tool as it defines states in the thermodynamic limit directly, avoiding finite size extrapolations.
@@ II.1 Parton ansatze
where $\langle ij\rangle$ ($\langle\langle ik\rangle\rangle$) denotes nearest (next-nearest) neighbor bonds and $\sigma$ is the spin index, $\sigma=\{\uparrow,\downarrow\}$. The hoppings are identical for both spin species. We fix $t_{2}=0.5t_{1}$, $\chi_{ij}=\pm 1$ to ensure a $\pi$ flux through every square plaquette and choose the complex phases $\theta_{ik}$ to obtain a $\pi/2$ flux in all triangles. In this paper we choose the gauge used in Ref. [66] such that the unit-cell can be chosen as two nearest neighbour sites along the $x$ direction.
The exact many-body spin wavefunction $|\varphi\rangle$ is a Gutzwiller projected Slater-determinant
@@
It is known that the resultant exact parton state is a $\rm SU(2)_{1}$ CSL which is equivalent to the $\nu=1/2$ bosonic Laughlin state. On a torus, such CSL has two-fold topological degeneracy where the degenerate states can be constructed by imposing different boundary conditions on the parton wavefunctions [67]. On a cylinder, the CSL hosts chiral gapless edge states predicted by $\rm SU(2)_{1}$ Wess-Zumino-Witten (WZW) CFT. In the rest of our work, we dub this parton state in (2) the ”exact CSL”, in contrast to its fPEPS approximation to be discussed below.
@@ III.2 Spin-spin correlations
Figure 2: Logarithmic plot of the magnitude of spin-spin correlations as a function of distance for the exact CSL and projected fPEPS, for the $L=18$ cluster with periodic boundary conditions. Symbols are ommited if the obtained value is lesser than one standard deviation of error. The expected correlations of the exact CSL at $r>5$ (in the thermodynamic limit) are shown as a dotted line, extending the exponential behavior.
@@ III.3 Topological properties: ground state degeneracy
$|\Psi^{\text{PBC-PBC}}\rangle&=\frac{|\Psi^{\text{PBC-APBC}}\rangle+|\Psi^{\text{APBC-PBC}}\rangle}{\sqrt{2}},\\ |\Psi^{\text{APBC-APBC}}\rangle&=\frac{|\Psi^{\text{PBC-APBC}}\rangle-|\Psi^{\text{APBC-PBC}}\rangle}{\sqrt{2}}\end{aligned}$ (8)
with a very good accuracy whenever $L\geq 18$, consistent with the fact that the states on the RHS and the LHS of the above equations belong to the same IRREPs of the $C_{4}$ rotation group on clusters with $L_{x}=L_{y}$ (respectively even and odd). These results substantiate that fPEPS, even at finite bond dimensions, can accurately capture the correct topological degeneracy of CSL states. This is one of the main results of our work.
@@ IV Simulation of the chiral $J_{1}-J_{2}-J_{\chi}$ model
This model was initially proposed in [63], where exact diagonalization techniques revealed that the ground state had a very high overlap with the exact Kalmeyer-Laughlin state. The model exhibits a chiral spin liquid (CSL) phase over a large region of its phase diagram, and we select a specific set of parameters also discussed in that work. The Hamiltonian can be undersdood as an effective Hamiltonian of the Fermi-Hubbard model through the Schrieffer–Wolff transformation. Realizing the Fermi-Hubbard model in cold atom experiments remains an active area of research [78, 79, 80]. Further, this model has been investigated in Ref. [54], where the sum of four triangular $J_{\chi}$ terms inside a plaquette has been equivalently written as a spin cyclic permutation $i(P_{ijkl}-P_{ijkl}^{-1})$ term. Ref. [54] performed a variational study of this model with bosonic iPEPS, and found a regime of $\rm SU(2)_{1}$ CSL in the phase diagram. The optimized bosonic iPEPS provides good variational energy and correct level counting in the ES predicted by $\rm SU(2)_{1}$ WZW CFT. However, there exists a redundant chiral branch in the odd (semion) sector [54, 81] which contradicts both the theoretical prediction and recent numerical results obtained from DMRG on finite cylinders [64, 65]. We emphasize that such artificial replication of chiral branches in bosonic iPEPS is quite general and is also found in the cases of $\rm SU(N)$ and non-Abelian CSLs [61, 62, 11, 55].
@@ IV.1 Variational energy
We choose the parameters of the spin model as $J_{1}=2\cos(0.06\pi)\cos(0.14\pi)$, $J_{2}=2\cos(0.06\pi)\sin(0.14\pi)$, $J_{\chi}=4\sin(0.06\pi)$ which has been considered in Refs. [63, 54, 76] and is known to be deep inside the CSL phase with the ground state energy being $E\approx-1$. We import the $\rm SU(2)$ bosonic iPEPS method there and compute variational energies as a reference. For the bosonic iPEPS, the tensor has $\rm SU(2)$ symmetry and the unit-cell size is chosen to be one such that each tensor only has one physical leg. The results for virtual spaces $V=0\oplus 1/2$ ($D=3$) and $V=0\oplus 0\oplus 1/2$ ($D=4$) are given in Fig. 4. As we extrapolate to the infinite $\chi$ limit the energies are around $-0.99$.
@@ V Conclusion and outlook
Tensor Network methods have been widely applied to study frustrated spin models hosting possible spin liquid phases. However, due to a no-go theorem [57], doubts have arisen regarding their ability to describe chiral spin liquid states. In this work, we investigated the topological properties of projected fPEPS ansatz for a CSL state. By performing a VMC analysis on the initial parton state, we demonstrated that fPEPS at finite bond dimension can accurately capture the topological GS degeneracy of CSLs. This achievement has been challenging with conventional bosonic PEPS due to the high computational cost involved [52]. Additionally, our results show that the fPEPS ansatz effectively captures the spin-spin correlation functions of the CSL. Given the high quality of overlap fidelity, we anticipate that fPEPS will also accurately reproduce the modular S and T matrices.
@@
Our work provides further supporting evidence for using PEPS methods in describing topologically ordered states. For example, fPEPS ansatz can further be used to study fermionic Hofstadter-Hubbard model [83] where both the Chern insulating phase and CSLs can be tuned by varying the strength of Hubbard interaction (e.g. Ref. [42]). In that case, our Gaussian fPEPS ansatze with partially projected doublons are expected to be good initial variational ansatze at finite Hubbard interaction.